OpenStudy (anonymous):

Can someone do the following question AND SHOW ALL WORK! Complete the square: y=x^2-11x+1

OpenStudy (jhannybean):

When you're completing the square, you're trying to attain a new c-value that will turn your given equation into a quadratic equation that is factorable.

OpenStudy (jhannybean):

\[f(x) = x^2 -11x +1\] group the x terms together and creae a quadratic inside the parenthesis by : \(c = (b/2)^2\)\[f(x) = (x^2 -11x) +1\]\[f(x) = \left(x^2 -11x +\frac{121}{4}\right) + 1 -\frac{121}{4}\]?Because we want to keep the intact, we have to add and subtract out values in a manner that will not change our original function. That is why if we found our new \(c\) value, we must also subtract by it accordingly. Now simplify. \[f(x) = \left(x-\frac{11}{2}\right)^2 -\frac{117}{4}\] That is your factored form using completing the square :)

OpenStudy (nincompoop):

pfftt, only a sucker would "show all work" so you can copy it and claim you did it.

OpenStudy (jhannybean):

Lmao -.-

OpenStudy (nincompoop):

i would rather show the concept to the fool and let him figure his life out early

OpenStudy (anonymous):

Not if they add this at the bottom \[\tiny Jhannybean\]

OpenStudy (jhannybean):

How did you do that!

OpenStudy (nincompoop):

they should include a citation that the solution were provided by somebody else

OpenStudy (anonymous):

\tiny Anyone know how to make piecewise functions

OpenStudy (nincompoop):

stop plagiarism and cheating

OpenStudy (anonymous):

LMFAO!! My favorite type of question, the one that basically screams " I'm cheating on my homework or test "